# DESI results and Dark Energy from QCD topological sectors

### v1

@@ III DESI results and deviation from exact de Sitter behaviour
The DESI results [3] suggest not only that the dark energy equation-of-state parameter $w$ differs from $-1$, but also that it may vary over time and even cross the $w=-1$ boundary. If confirmed, such behavior would present significant challenges for conventional field theory models. The aim of this section is to examine whether this kind of behavior can arise within our proposed framework. To do so, we begin by representing the Friedmann equation in the context of our approach and analyzing its solutions.
@@ III.2 The de Sitter limit
One can clearly see that $w\rightarrow-1$ as $x(t)\rightarrow 0$, recovering our earlier Eq. (7) for the de Sitter state. Interestingly, as we approach the present time, when $x(t_{0})\sim 1$, Eq.(18) yields $w\simeq-0.7$, illustrating two important consequences of our framework, in agreement with the DESI results: (1) the equation-of-state parameter $w$ is time-dependent, and (2) it is indeed possible to have $w>-1$ at the present epoch.
@@ III.3 Solution for $z\geq 0$
The quantity $\beta(t)$ functions as a switch that activates dark energy at a certain redshift. It can be interpreted as a physically motivated parametrization of dark energy within our framework, where dark energy is induced by QCD. This contrasts with the canonical parametrization $w(a)=w_{0}+w_{a}(1-a)$, which merely captures a linear deviation from $w=-1$. It is important to emphasize that introducing a time-dependent $\beta(t)$ is not an ad hoc addition to the proposal. Rather, it serves as an effective phenomenological tool to constrain the region of validity of our formulae near the de Sitter regime, where they were derived—hence its physical motivation. In principle, as previously mentioned, the numerical coefficient $[\beta(t)c_{H}]$ can be calculated from first principles for any given geometry at a given time $t$, followed by the appropriate subtraction of the corresponding Minkowski space value, in analogy with the subtraction procedure performed in (II.3) for pure de Sitter space with constant $\overline{H}$.
@@
In practice, the solution for $H(a)$ derived in Section III.1 remains valid, with the exception that $[\beta(t)c_{H}]$ is now explicitly time-dependent. Solving for $a(t)$ from Eq. (11) can only be done numerically, provided that $\beta(t)$ is specified. Consequently, our framework will modify the solution to the Friedmann equation, Eq. (11), modification which must be tested against all standard cosmological observations  [^6] All cosmological constraints, from the cosmic microwave background, baryon accoustic oscillations, supernovae and large scale structures tests will have to be reprocessed with this new framework to assess its validity. In this context, $\beta(t)$ can be treated as a parameterized function related to DE, although it is not the DE equation of state itself. This represents a substantial undertaking and is left for future work.
@@
What can be done at this stage is to illustrate the impact of a time-dependent $\beta(t)$ on cosmology, in comparison to the standard $\Lambda$CDM model. To this end, we will consider two different forms of $\beta(t)$: one that activates dark energy gradually over cosmic time, and another that activates it almost like a step function. In order to ensure a fair comparison between these scenarios and the conventional $\Lambda$CDM cosmology, we will solve Eq. (11) using an anchor time $i$ set sufficiently far in the past, when dark energy was negligible and only radiation and matter contributed significantly. It is convenient to redefine time $t$ as $\tau\equiv(\frac{8\pi G}{3}\Lambda_{\rm QCD}^{3})c_{H}t$ and rewrite Eq.(11) as:
@@
Equation (22) is then solved numerically with $\overline{H}=50\;{\rm km\;s^{-1}\;Mpc^{-1}}$. Figure 2 shows the variation of the Hubble parameter $H(z)$, the dark energy density $\rho_{\rm DE}(z)$, the density parameter $\Omega_{\rm DE}(z)$, and the dark energy equation of state $w_{\rm QCD}(z)$ as functions of redshift, and compares them to the predictions of the Planck cosmology. Panels (a), (b), and (c) illustrate that different histories of QCD dark energy activation can influence key cosmological quantities, which are fundamental for cosmological tests. Panel (a) demonstrates that a late-time change in the Hubble constant is possible, potentially indicating a connection with the $H_{0}$ tension, as discussed in Section III.4.
Panel (d) shows that, within our framework, the dark energy equation of state parameter can vary with time, and may even cross the $w=-1$ line, even multiple times. Although our choice of the $\beta(t)$ function is entirely arbitrary and not based on any physical model, it is still possible to test the self-consistency of the framework by performing all relevant cosmological tests. If a single $\beta(t)$ function proves successful across all these tests, it would provide strong evidence in support of our approach. The main result of this numerical experiment is that it is straightforward to obtain an equation of state parameter that varies with time and can take values both above and below $-1$.
Panel (d) also shows that for $z<0$ (i.e., in the future), $w_{\rm DE}$ converges to $-1$, consistent with the derivation in Section III.2. This is a generic feature of our framework: we predict that the de Sitter phase will be asymptotically reached in the future, implying that today we should observe $w_{\rm DE}>-1$. The commonly used parametrization $w(a)=w_{0}+w_{a}(1-a)$ does not represent a good model in our framework because it does not capture the phase towards de Sitter described in Section III.2. For $z>0$, however, $w_{\rm DE}$ can take any form as a function of redshift, depending on the activation function of the QCD-induced dark energy, $\beta(t)$. Although we cannot derive $\beta(t)$ theoretically, it is likely to be strongly constrained by observational data. Moreover, by comparing panels (a) and (d), one can see a strong correlation between the Hubble parameter $H(z)$ and $w(z)$. This correlation is unique to our framework because it is coming from the fact that $\rho_{\rm DE}(t)\propto\beta(t)H(t)$ in recent times when DE started to dominate the evolution of the universe. It offers an additional testable prediction of the framework. What is needed for this test are independent measurements of both $H(z)$ and $w(z)$. Future data releases from DESI and Stage IV surveys (such as Euclid and LSST) will improve the precision of $w(z)$. While current independent measurements of the Hubble parameter at high redshifts are still imprecise, there is hope for significant improvement in the near future.
@@ III.4 Possible relation to the $H_{0}$ tension?
It is well known that the canonical approach, in which $\rho_{\rm DE}$ is treated as a cosmological constant, may be related to several observed inconsistencies, as discussed, for example, in [22], or a recent observation that unphysical negative neutrino mass reduces some tension [23]. Could some of these inconsistencies be resolved if $\rho_{\rm DE}\propto H$, as proposed in this work? In particular, might the $H_{0}$ tension be one such case? By comparing the present-day value of $H_{0}$ derived within the canonical $\Lambda$CDM framework to that obtained in our approach, one can evaluate the extent to which the two Hubble constants may differ.
We present below a simple analytical argument suggesting that this difference is, in fact, a generic consequence of the framework proposed in this work. Indeed, for the QCD-induced dark energy, Eq.(12) applies, and we can denote $H_{0}$ as $H_{0}({\rm QCD})$ to clearly associate it with our framework. By squaring Eq.(12), we obtain the following algebraic relation:
$\frac{H_{0}({\rm QCD})-\overline{H}}{H_{0}({\rm QCD})}=\Omega_{M,0},~{}\rightarrow~{}\frac{H_{0}({\rm QCD})}{\overline{H}}=\frac{1}{(1-\Omega_{M,0})}.~{}~{}$ (25)
We can now proceed with similar calculations under the assumption of the canonical $\Lambda$CDM scenario. In this case, Eq.(10) applies, with the dark energy term $\overline{H}H$ replaced by $\overline{H}^{2}$, since in $\Lambda$CDM the dark energy component is assumed to be constant. We identify this constant with the asymptotic value determined by our framework. At the present time ($i=0$), Eq.(10) can be rewritten as:
$\frac{H_{0}(\Lambda CDM)}{\overline{H}}=\frac{1}{\sqrt{(1-\Omega_{M,0})}},$ (26)
where $H_{0}(\Lambda CDM)$ is the Hubble constant in the canonical Friedmann case.
It is important to emphasize that the same value of $\overline{H}$ should be used in both cases, as it serves as the normalization point for $H$ in the future, where the universe is 100% dominated by the de Sitter state. By comparing Eqs. (25) and (26), we arrive at the conclusion that:
$\frac{H_{0}(\rm QCD)}{H_{0}(\Lambda CDM)}=\frac{1}{\sqrt{(1-\Omega_{M,0})}}>1.$ (27)
Eq.(27) shows that the value of $H_{0}$ measured today, $H_{0}({\rm QCD})$, is greater than the value predicted by a canonical $\Lambda CDM$ Friedmann cosmology, $H_{0}({\Lambda CDM})$, such as that inferred from CMB measurements extrapolated to $z=0$.
This suggests a possible link between our proposal and the $H_{0}$ tension. Our simple analytical arguments are supported by numerical computations from Sect. III.3. Indeed, panel (a) of Figure 2 illustrates that for both $\beta(t)$ cases considered here, the resulting Hubble constant today is larger than the value extrapolated from Planck data, which is consistent with our analytical argument (27). A definitive test of this hypothesis will require a full MCMC analysis of our framework against cosmological observations since distances and time scales might be modified in our framework [24].
@@ IV Conclusion
We have presented a physically motivated model of dark energy (DE) rooted in the topological structure of the QCD vacuum, as developed in earlier works [7, 8, 9]. In this framework, DE is identified with the leftover vacuum energy generated by tunnelling transitions between different topological sectors, following the subtraction of a large contribution computed in Minkowski spacetime—an approach known as the Zeldovich prescription [10]. The amplitude of DE is determined by the QCD energy scale, as given in Eq. (3), while its time dependence arises from the evolution of the topological sectors $|k\rangle$ and the corresponding variation in tunnelling transition rates in an expanding universe.
@@
The key new element introduced in this work is the extension of earlier ideas [7, 8, 9] to account for deviations from a pure de Sitter state. This study is fully motivated by recent DESI results [3], which clearly indicate significant departures from the standard $\Lambda$CDM model. While additional dark energy measurements from independent experiments are needed to confirm these findings, we demonstrate that such deviations are not unexpected within our framework. Moreover, we propose a variety of cosmological tests to challenge and validate our approach. Our main findings are as follows:
@@
2.For $z\geq 0$, the precise value and redshift dependence of $w_{\rm DE}(z)$ are governed by an activation function $\beta(t)$, which controls the emergence and growth of QCD-induced dark energy over cosmic time in close vicinity of the de Sitter state. While this function cannot be computed analytically, it influences all cosmological observables—such as timescales and distance measures—and thereby affects standard candles and rulers. This makes the framework empirically testable with current cosmological data, though a comprehensive analysis is left for future work. If any observational deviation from $\Lambda$CDM—consistent with the predictions of our framework—is confirmed, it would constitute strong evidence in support of our model.
3. The QCD-induced dark energy leads to solutions of the Friedmann equation in which $w_{\rm DE}$ may lie above or below $-1$ and can cross the $w=-1$ boundary multiple times throughout the history of the Universe, a trend suggested by the DESI results [3].
@@
5. The de Sitter regime described by the equation of state (7) would represent the final state of our Universe if the interaction between QCD gauge configurations (which saturate the vacuum energy) and massless electromagnetic photons were permanently switched off, as noted in item 1 above. However, when the coupling between the QCD vacuum fields and the electromagnetic field is restored, the departure from de Sitter behavior is triggered precisely by this interaction, which is unambiguously determined by the triangle anomaly, as discussed in [25].
@@
6. Finally, the QCD-induced dark energy framework predicts the generation of a cosmic magnetic field with an estimated strength on the order of $10^{-10}$ G, as argued in [25]. Remarkably, the corresponding correlation length spans the scale of the observable Universe—on the order of gigaparsecs. Intriguingly, observational evidence for magnetic fields with such enormous correlation lengths has indeed been reported, as reviewed in [26].
@@
Interestingly, is it, at least in principle, possible to test some of the key ideas underlying this dark energy framework—specifically its origin in tunnelling processes between different topological sectors—using tabletop experiments? The ultimate answer is affirmative: this represents a genuine physical phenomenon rather than a mere formal reinterpretation of equations. The basic concept behind such an experiment is to detect a novel contribution to the Casimir vacuum energy in Maxwell theory, as proposed in [27, 28, 29]. This contribution, known as the Topological Casimir Effect, arises from non-dispersive tunnelling effects rather than from conventional quantum fluctuations of propagating photons with two physical transverse polarizations (dispersive contribution). Although this correction to the Casimir pressure has not yet been observed, its detection would provide direct empirical support for the topological origin of dark energy.
@@
Our final remark concerning possible future developments of this work is as follows. It is well known that de Sitter–like behavior has occurred twice in the history of the Universe: first during the inflationary epoch, and again in the present epoch dominated by dark energy. The dark energy framework explored in this paper may offer valuable insights into the inflationary phase. Indeed, in a purely hypothetical scenario proposed in [7, 9], the vacuum energy responsible for inflation could also arise from tunnelling transitions in a novel, as yet unidentified, strongly coupled gauge theory—an idea analogous to the QCD-based mechanism we advocate here for generating the dark energy scale, as described in Eq. (3). While the specific gauge theory that might play the role of QCD in this inflationary context remains unknown, the relevant energy scale (analogous to $\Lambda_{\mathrm{QCD}}$ for dark energy) and the expected number of e-foldings can nevertheless be estimated within this framework [7, 9].
@@ Appendix A On use and misuse of the so-called Ghost DE model
The underlying physics can indeed be described using auxiliary topological fields, which offer an alternative to explicit non-perturbative calculations, as reviewed in Section II.2. These fields are non-dynamical, as they lack a canonically conjugate momentum. A concrete construction of such an auxiliary topological field has been carried out in the exactly solvable deformed QCD model [21], where essential features such as topological sectors $|k\rangle$ and non-trivial holonomy are explicitly realized. In this context, it becomes evident that the auxiliary field introduced in [21] does not propagate and possesses no kinetic term. Instead, it functions purely as a Lagrange multiplier within the system. Moreover, the energy described by Eq. (6) and generated by this auxiliary field cannot be attributed to any propagating degrees of freedom—it constitutes a so-called non-dispersive contribution.
It turns out that these topological fields exhibit precisely the properties of the Veneziano ghost [15], which was originally postulated to resolve the so-called $U(1)_{A}$ problem in QCD.  [^7] To our knowledge, the explicit construction of the Veneziano ghost in an exactly solvable model, as presented in [21], remains the only known example where one can observe the development of a $1/q^{2}$ pole at zero momentum with a residue of the “wrong” sign—hence the designation as a ghost. The same construction also demonstrates that the Veneziano ghost is non-propagating, thereby avoiding any conflict with fundamental principles of quantum field theory such as unitarity and causality. This connection between dark energy and the Veneziano ghost, as originally proposed in [19, 20], led to the model being named “Ghost Dark Energy” (GDE). A large body of literature has since explored various aspects and generalizations of this idea. However, many of these works have been accompanied by significant misconceptions regarding the theoretical foundations of GDE.
@@
In particular, in [30] and many subsequent papers, the expression given in Eq. (6) was interpreted as if it were generated by a dynamical field $\phi$, with the equation of state (EoS) and the speed of sound $c_{s}$ expressed in terms of this field $\phi$ as follows:
$w=\frac{\dot{\phi}^{2}-2V(\phi)}{\dot{\phi}^{2}+2V(\phi)},~{}~{}~{}~{}~{}~{}~{}c_{s}^{2}\equiv\frac{{\dot{p}_{DE}}}{{\dot{\rho}_{DE}}}.$ (28)
From the expressions for $w$ and $c_{s}$, it becomes evident that the system exhibits serious issues—such as violations of unitarity, causality, and stability—when $w<-1$ or $c_{s}^{2}<0$. However, these problems arise only when the underlying model assumes a dynamical field. In contrast, when the variation originates from an auxiliary (non-dynamical) field, as in our framework, such behavior does not conflict with any fundamental principles, since there are no propagating degrees of freedom—there is simply nothing to propagate. Consequently, concerns about classical stability for cases where $c_{s}^{2}<0$, as discussed for example in [31, 32], are not relevant to this model. The computation of a speed of sound in this context is meaningless, as there is no physical field to support fluctuations.
Another common misconception is as follows. If one interprets the behavior with $w<-1$—which is precisely what we have found in Section III—as indicative of phantom dark energy, then, following Eq. (28), one could reconstruct a specific scalar potential $V(\phi)$ that would yield such an equation of state, as discussed in [33]. However, as we have emphasized, there are no propagating degrees of freedom in our framework. The occurrence of $w<-1$ is entirely consistent with the fundamental principles of quantum field theory, precisely because it does not involve any dynamical field. Consequently, within our model, describing a regime with $w<-1$ does not necessitate the introduction of a fundamental phantom field, in stark contrast to the conventional interpretation.
To conclude this Appendix, which has addressed common misinterpretations of Eq. (6) and its various implications—including the regime with $w<-1$—we emphasize that the so-called phantom behavior in our framework is fully consistent with the fundamental principles of quantum field theory. This consistency arises from the fact that the underlying dynamics is not associated with any propagating or dynamical degrees of freedom.

### v2

@@ III DESI results and deviation from exact de Sitter behaviour
The DESI results DESI Collaboration et al. (2025) suggest not only that the dark energy equation-of-state parameter $w$ differs from 1-1, but also that it may vary over time and even cross the w=1w=-1 boundary. If confirmed, such behavior would present significant challenges for conventional field theory models. The aim of this section is to examine whether this kind of behavior can arise within our proposed framework. To do so, we begin by representing the Friedmann equation in the context of our approach and analyzing its solutions.
@@ III.2 The de Sitter limit
One can clearly see that $w\rightarrow-1$ as $x(t)\rightarrow 0$, recovering our earlier Eq. (7) for the de Sitter state. Interestingly, as we approach the present time, when $x(t_{0})\sim 1$, Eq.(18) yields $w\simeq-0.7$, illustrating two important consequences of our framework, in agreement with the DESI results: (1) the equation-of-state parameter $w$ is time-dependent, and (2) it is indeed possible to have w>1w>-1 at the present epoch.
@@ III.3 Solution for $z\geq 0$
The quantity $\beta(t)$ functions as a switch that activates dark energy at a certain redshift. It can be interpreted as a physically motivated parametrization of dark energy within our framework, where dark energy is induced by QCD. This contrasts with the canonical parametrization $w(a)=w_{0}+w_{a}(1-a)$, which merely captures a linear deviation from w=1w=-1. It is important to emphasize that introducing a time-dependent $\beta(t)$ is not an ad hoc addition to the proposal. Rather, it serves as an effective phenomenological tool to constrain the region of validity of our formulae near the de Sitter regime, where they were derived—hence its physical motivation. In principle, as previously mentioned, the numerical coefficient $\beta(t)c_{H}]$ can be calculated from first principles for any given geometry at a given time $t$, followed by the appropriate subtraction of the corresponding Minkowski space value, in analogy with the subtraction procedure performed in (II.3) for pure de Sitter space with constant $\overline{H}$.
@@
[…]
@@
[…]
@@
Equation (22) is then solved numerically with $\overline{H}=50\;{\rm km\;s^{-1}\;Mpc^{-1}}$. Figure 2 shows the variation of the Hubble parameter $H(z)$, the dark energy density $\rho_{\rm DE}(z)$, the density parameter $\Omega_{\rm DE}(z)$, and the dark energy equation of state $w_{\rm QCD}(z)$ as functions of redshift, and compares them to the predictions of the Planck cosmology. Panels (a), (b), and (c) illustrate that different histories of QCD dark energy activation can influence key cosmological quantities, which are fundamental for cosmological tests. Panel (d) shows that, within our framework, the dark energy equation of state parameter can vary with time, and may even cross the w=1w=-1 line, even multiple times. Although our choice of the $\beta(t)$ function is entirely arbitrary and not based on any physical model, it is still possible to test the self-consistency of the framework by performing all relevant cosmological tests. If a single $\beta(t)$ function proves successful across all these tests, it would provide strong evidence in support of our approach. The main result of this numerical experiment is that it is straightforward to obtain an equation of state parameter that varies with time and can take values both above and below 1-1 without the need for a new dynamical field.
Panel (d) also shows that for $z<0$ (i.e., in the future), $w_{\rm DE}$ converges to 1-1, consistent with the derivation in Section III.2. This is a generic feature of our framework: we predict that the de Sitter phase will be asymptotically reached in the future, implying that today we should observe $w_{\rm DE}>-1$. The commonly used parametrization $w(a)=w_{0}+w_{a}(1-a)$ does not represent a good model in our framework because it does not capture the phase towards de Sitter described in Section III.2. For $z>0$, however, $w_{\rm DE}$ can take any form as a function of redshift, depending on the activation function of the QCD-induced dark energy, $\beta(t)$. Although we cannot derive $\beta(t)$ theoretically, it is likely to be strongly constrained by observational data. Moreover, by comparing panels (a) and (d), one can see a strong correlation between the Hubble parameter $H(z)$ and $w(z)$. This correlation is unique to our framework because it is coming from the fact that $\rho_{\rm DE}(t)\propto\beta(t)H(t)$ in recent times when DE started to dominate the evolution of the universe. It offers an additional testable prediction of the framework. What is needed for this test are independent measurements of both $H(z)$ and $w(z)$. Future data releases from DESI and Stage IV surveys (such as Euclid and LSST) will improve the precision of $w(z)$. While current independent measurements of the Hubble parameter at high redshifts are still imprecise, there is hope for significant improvement in the near future.
@@ III.4 Possible relation to the $H_{0}$ tension?
Figure 3: Commoving distance as a function redshift $z$ relative to the Planck cosmology. The red abd black curves correspond to the activation functions $\beta_{1}(t)$ and $\beta_{2}(t)$ in Eq.(24).
[…]
$\theta_{\star}\equiv\frac{r_{s}(z_{\star})}{D_{A}(z_{\star})},$ (25)
where $D_{A}(z_{\star})$ is the angular diameter distance at recombination and $r_{s}(z_{\star})$ is the sound horizon at $z_{\star}$. It is determined by the early-time integral of the sound speed $c_{s}(z)$:
$r_{s}(z_{\star})=\int_{z_{\star}}^{\infty}\frac{c_{s}(z)}{H(z)}dz$ (26)
On the other hand, BAO measurement take place at some late-time redshift $z_{\rm low}$. It measures the sound horizon angular scale $\theta_{d}(z_{\rm low})$:
$\theta_{d}\equiv\frac{r_{s}(z_{d})}{D_{A}(z_{\rm low})},$ (27)
with $r_{s}(z_{d})\simeq 1.018r_{s}(z_{\star})$ and where $z_{d}\sim 1059$ is the baryong drag redshift, which takes place slightly after recombination $z_{\star}\sim 1090$, when baryons are fully released from the photons drag. The impossibility of resolving the Hubble tension through a single early- or late-time modification arises from the fact that Eqs. (25) and (27) cannot be simultaneously satisfied under current cosmological constraints (Bernal et al., 2016).
It has been suggested that combining pre- and post-recombination new physics may offer a more satisfactory resolution of the $H_{0}$ tension, while simultaneously alleviating the amplitude-of-fluctuations tension in $S_{8}$ (Vagnozzi et al., 2022; Adil et al., 2023; Poulin et al., 2024). In our framework, the Friedmann equation is modified at all redshifts, with the detailed behavior determined by the unknown function $\beta(z)$. Using the two ad hoc choices for $\beta(z)$ introduced in Eq. (24), Figure 3 demonstrates that comoving distances can be altered across the entire redshift range. This behavior parallels the evolution of the Hubble constant shown in Panel (a) of Fig. 2. Therefore, an important open question is whether there exists a function $\beta(z)$ capable of satisfying all current cosmological constraints from CMB, BAO, SN, and LSS, how many free parameters such a function would require, and to what extent it could provide a physically viable description of the QCD vacuum. A full statistical analysis using a realistic parametrization of $\beta(z)$ is left to future work.
Note that, under the conjecture that Eq. (1) applies to the Hubble parameter, spatial variations in the topological sectors of the QCD vacuum may arise because local energy-density fluctuations would modify the Hubble rate. This opens the possibility of behavior resembling clustered dark energy, or more precisely, a dark-energy equation of state that correlates with the local mass density. Within our framework, such behavior can emerge without introducing any exotic additional fields, and may provide a new avenue for dark energy coupled to the dark sector (Pitrou and Uzan, 2024).
@@ IV Conclusion
We have presented a physically motivated model of dark energy (DE) rooted in the topological structure of the QCD vacuum, as developed in earlier works Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018). In this framework, DE is identified with the leftover vacuum energy generated by tunnelling transitions between different topological sectors, following the subtraction of a large contribution computed in Minkowski spacetime—an approach known as the Zeldovich prescription Zeldovich (1967). The amplitude of DE is determined by the QCD energy scale, as given in Eq. (3), while its time dependence arises from the evolution of the topological sectors $|k\rangle$ and the corresponding variation in tunnelling transition rates in an expanding universe.
@@
[…]
@@
[…]
3. The QCD-induced dark energy leads to solutions of the Friedmann equation in which $w_{\rm DE}$ may lie above or below 1-1 and can cross the w=1w=-1 boundary multiple times throughout the history of the Universe, a trend suggested by the DESI results DESI Collaboration et al. (2025).
@@
5. The de Sitter regime described by the equation of state (7) would represent the final state of our Universe if the interaction between QCD gauge configurations (which saturate the vacuum energy) and massless electromagnetic photons were permanently switched off, as noted in item 1 above. However, when the coupling between the QCD vacuum fields and the electromagnetic field is restored, the departure from de Sitter behavior is triggered precisely by this interaction, which is unambiguously determined by the triangle anomaly, as discussed in Zhitnitsky (2019).
@@
6. Finally, the QCD-induced dark energy framework predicts the generation of a cosmic magnetic field with an estimated strength on the order of $10^{-10}$ G, as argued in Zhitnitsky (2019). Remarkably, the corresponding correlation length spans the scale of the observable Universe—on the order of gigaparsecs. Intriguingly, observational evidence for magnetic fields with such enormous correlation lengths has indeed been reported, as reviewed in Durrer and Neronov (2013).
@@
Interestingly, is it, at least in principle, possible to test some of the key ideas underlying this dark energy framework—specifically its origin in tunnelling processes between different topological sectors—using tabletop experiments? The ultimate answer is affirmative: this represents a genuine physical phenomenon rather than a mere formal reinterpretation of equations. The basic concept behind such an experiment is to detect a novel contribution to the Casimir vacuum energy in Maxwell theory, as proposed in Cao et al. (2013, 2016); Yao and Zhitnitsky (2017). This contribution, known as the Topological Casimir Effect, arises from non-dispersive tunnelling effects rather than from conventional quantum fluctuations of propagating photons with two physical transverse polarizations (dispersive contribution). Although this correction to the Casimir pressure has not yet been observed, its detection would provide direct empirical support for the topological origin of dark energy.
@@
Our final remark concerning possible future developments of this work is as follows. It is well known that de Sitter–like behavior has occurred twice in the history of the Universe: first during the inflationary epoch, and again in the present epoch dominated by dark energy. The dark energy framework explored in this paper may offer valuable insights into the inflationary phase. Indeed, in a purely hypothetical scenario proposed in Zhitnitsky (2014); Barvinsky and Zhitnitsky (2018), the vacuum energy responsible for inflation could also arise from tunnelling transitions in a novel, as yet unidentified, strongly coupled gauge theory—an idea analogous to the QCD-based mechanism we advocate here for generating the dark energy scale, as described in Eq. (3). While the specific gauge theory that might play the role of QCD in this inflationary context remains unknown, the relevant energy scale (analogous to $\Lambda_{\mathrm{QCD}}$ for dark energy) and the expected number of e-foldings can nevertheless be estimated within this framework Zhitnitsky (2014); Barvinsky and Zhitnitsky (2018).
@@ Appendix A On use and misuse of the so-called Ghost DE model
The underlying physics can indeed be described using auxiliary topological fields, which offer an alternative to explicit non-perturbative calculations, as reviewed in Section II.2. These fields are non-dynamical, as they lack a canonically conjugate momentum. A concrete construction of such an auxiliary topological field has been carried out in the exactly solvable deformed QCD model Zhitnitsky (2013), where essential features such as topological sectors $|k\rangle$ and non-trivial holonomy are explicitly realized. In this context, it becomes evident that the auxiliary field introduced in Zhitnitsky (2013) does not propagate and possesses no kinetic term. Instead, it functions purely as a Lagrange multiplier within the system. Moreover, the energy described by Eq. (6) and generated by this auxiliary field cannot be attributed to any propagating degrees of freedom—it constitutes a so-called non-dispersive contribution.
It turns out that these topological fields exhibit precisely the properties of the Veneziano ghost Veneziano (1979), which was originally postulated to resolve the so-called $U(1)_{A}$ problem in QCD.  [^7] To our knowledge, the explicit construction of the Veneziano ghost in an exactly solvable model, as presented in Zhitnitsky (2013), remains the only known example where one can observe the development of a $1/q^{2}$ pole at zero momentum with a residue of the “wrong” sign—hence the designation as a ghost. The same construction also demonstrates that the Veneziano ghost is non-propagating, thereby avoiding any conflict with fundamental principles of quantum field theory such as unitarity and causality. This connection between dark energy and the Veneziano ghost, as originally proposed in Urban and Zhitnitsky (2010a, b), led to the model being named “Ghost Dark Energy” (GDE). A large body of literature has since explored various aspects and generalizations of this idea. However, many of these works have been accompanied by significant misconceptions regarding the theoretical foundations of GDE.
@@
In particular, in Cai et al. (2011) and many subsequent papers, the expression given in Eq. (6) was interpreted as if it were generated by a dynamical field $\phi$, with the equation of state (EoS) and the speed of sound $c_{s}$ expressed in terms of this field $\phi$ as follows:
$w=\frac{\dot{\phi}^{2}-2V(\phi)}{\dot{\phi}^{2}+2V(\phi)},~~~~~~~c_{s}^{2}\equiv\frac{{\dot{p}_{DE}}}{{\dot{\rho}_{DE}}}.$ (A1)
From the expressions for $w$ and $c_{s}$, it becomes evident that the system exhibits serious issues—such as violations of unitarity, causality, and stability—when w<1w<-1 or $c_{s}^{2}<0$. However, these problems arise only when the underlying model assumes a dynamical field. In contrast, when the variation originates from an auxiliary (non-dynamical) field, as in our framework, such behavior does not conflict with any fundamental principles, since there are no propagating degrees of freedom—there is simply nothing to propagate. Consequently, concerns about classical stability for cases where $c_{s}^{2}<0$, as discussed for example in Ebrahimi and Sheykhi (2011); Biswas et al. (2019), are not relevant to this model. The computation of a speed of sound in this context is meaningless, as there is no physical field to support fluctuations.
Another common misconception is as follows. If one interprets the behavior with w<1w<-1—which is precisely what we have found in Section III—as indicative of phantom dark energy, then, following Eq. (A1), one could reconstruct a specific scalar potential $V(\phi)$ that would yield such an equation of state, as discussed in Das and Bharali (2021). However, as we have emphasized, there are no propagating degrees of freedom in our framework. The occurrence of w<1w<-1 is entirely consistent with the fundamental principles of quantum field theory, precisely because it does not involve any dynamical field. Consequently, within our model, describing a regime with w<1w<-1 does not necessitate the introduction of a fundamental phantom field, in stark contrast to the conventional interpretation.
To conclude this Appendix, which has addressed common misinterpretations of Eq. (6) and its various implications—including the regime with w<1w<-1—we emphasize that the so-called phantom behavior in our framework is fully consistent with the fundamental principles of quantum field theory. This consistency arises from the fact that the underlying dynamics is not associated with any propagating or dynamical degrees of freedom.